If your salary is reduced by 25%, by what percent would it then have to be increased to bring it back to the original amount?
Answer 33 1/3% (How?)

Found 2 solutions by humzaali, Theo:
Answer by humzaali(1)

Suppose innitially you have 100 rupees . your salary then decreased by 25% so it became 75 (25% of 100 is 25). Now you want to raise your salary again to 100 so you need increase of 25 in your salary
So we will use this condition to find the percent change
25 = x/100 × 75 whare x is required percentage change
And x will eqaul 33.333..

Answer by Theo(13342)


assume the original value is 100.

25% of 100 is equal to 1/4 * 100 which is equal to 25.

(33 + 1/3)% of 75 is equal to 1/3 * 75 which is equal to 25.

the difference is the same but the percent difference is not because the denominator in the equation is not the same.

25/100 = .25

25/75 = .3333.....

when you subtract 25% of 100 from 100, you are subtracting 25.

when you add 33 and 1/3% of 75 to 75, you are adding 25.

that's about as simple as i can explain it.

if you have further questions, let me know.